# PLAICraft: Large-Scale Time-Aligned Vision-Speech-Action Dataset for Embodied AI

### v1

@@ 4 Evaluation
| Test Type | Prompt | Response | Metric |
| audio to action binding | Press A. | A pressed. | |
@@
| spelling ability | Please press the letter that ends the word STOP, as in S-T-O. What letter was that? | P pressed. | |
| object recognition | What is this? What am I fighting? | Zombie. | |
| world-model understanding | Right now, will the spider attack us? | No. | |
@@
| arithmetic | If you add up the blocks in the wool and granite pile, how many are there? | Five. | |
| long-term memory | Have you seen this chest before? | Yes. | |
| object constancy | Jump one time. Can you see me? | Yes. | |
| spatial understanding | How many blocks away is the box? | Hundred. |
@@
Motivated by Animal-AI (Voudouris et al., 2022), which adopts testing paradigms from comparative psychology; BIB (Gandhi et al., 2022), which assesses machines’ abilities to reason about other agents’ intentions by observing their actions; and the taxonomy of cognitive abilities outlined in Cattell-Horn-Carroll (CHC) theory (Carroll, 1993), we publish an evaluation suite that probes how agents perceive, understand, reason, and act within PLAICraft. The idea behind this design is intuitive: if an agent truly understands what it sees and hears, it should be able to (i) talk about it, (ii) act on it, and (iii) do so within a time-bounded, social context.
@@
Our evaluation suite contains carefully curated prompts. Each prompt maps to one of the ten broad cognitive abilities in CHC theory (Schneider and McGrew, 2018): fluid intelligence, quantitative knowledge, crystallized intelligence, short-term memory, visual processing, auditory processing, long-term retrieval, processing speed, decision speed, and reading/writing. Each prompt elicits a keyboard/mouse action or a concise verbal answer, enabling evaluation via exact-match or semantic-similarity metrics.
@@
During evaluation, the agent’s prompt consists of its entire lived life up to the current moment, including every observation and interaction so far, and is then followed by a spoken instruction that triggers the task. A correct response from the agent may be a mouse & keyboard action ("A pressed") or a brief verbal answer ("Zombie"). The test type categorizes the cognitive ability being assessed, while the metric is the measure used to judge responses’ correctness.
Table 1 selects eight representative prompts from the evaluation suite in Appendix B. The tasks span from visual object recognition ("What am I fighting?"), long-term memory ("Have you seen this chest before?"), to spatial estimation ("How far is the box?") and arithmetic challenges. These tasks require agents to integrate visual, audio, and action in a temporally grounded context. We hope this focus on semantically rich evaluation will encourage research toward EAI agents that reason and play like real people.
5 Conclusion & Discussion
In this paper, we present a novel data-collection platform and accompanying dataset that together create new opportunities for research on open-ended learning, reinforcement learning, complex social interaction, memory consolidation, and other related topics.
We do acknowledge that our work has some limitations, for example, the social and audio interaction portion of the dataset is still relatively small, as we give players freedom to not interact with other players if they don’t want to. We have limited deployment regions due to cost considerations and have to focus on English-speaking regions. The evaluation suite is still not comprehensive enough and needs further exploration. Despite the complexity of the dataset, it is still not a real-world dataset.
But we are actively developing to address these limitations, and this is just our first step. Looking ahead, we are leveraging this dataset to train human-like embodied agents in Minecraft that can converse and cooperate with other players in real time. One of the examples, as demonstrated in Figure 8, Yoo et al. (2024) leveraged a 50-hour subset of our dataset—collected from an anonymous player (“Alex”)—to illustrate that a video-diffusion model can be effectively trained from a continuous gameplay stream in a continual learning setting (Wang et al., 2024). Moving away from Minecraft, because the platform is flexible and modality-agnostic, it can also be repurposed with minimal effort to collect multimodal data from any computer workflow, for example, screen-reading digital assistants, and eventually towards robotics.
@@ A.1 Acknowledgment
We are very grateful to Alexander Liteplo, Alex Wu, Andrew Smith, Andy Huang, Brooke Dai, Chengyuan Yao, Christopher Tardy, Daniel Crookall, David Tianyi Yin, David Yang, Edward Liang, Geo Lee, Hanson Sun, Joey Xiang, Lucas Qin, Mantaj Dhillon, Mehdi Safaee, Naveed Ghassemi, Perry Zhu, Sean Chuah, and Suzette Sun for their help in the development of the platform. We thank Alice Xia, Hotslicer Media, and many others for their promotion efforts. We also appreciate the help with server moderation and community organization by David Yang, Piper, Delara, and Hiroshi (the last three are anonymous players on the server who spontaneously offered to help with moderation). In addition, we would like to send a special thanks to all the amazing players who have spent countless hours on our server, enjoying Minecraft, building a wholesome community and contributing to our research.
@@
A.2 Related Work
Minecraft for Machine Learning Minecraft serves as an important tool for machine learning research [Lifshitz et al., 2024, Zhou et al., 2024] because of its open-world nature and intricate environment, along with a customizable game engine. Consequently, numerous datasets and benchmarks [Guss et al., 2019, Milani et al., 2023, Fan et al., 2022] have been established on Minecraft to enhance AI abilities. Microsoft’s first release of MALMO [Johnson et al., 2016] platform garnered great research interest in Minecraft. Based on MALMO, MineRL [Guss et al., 2019] introduces a complex environment for reinforcement learning tasks using human demonstrations in Minecraft. MineDojo [Fan et al., 2022] extends the capabilities by aggregating gameplay data from a variety of internet sources and providing a more expansive and flexible framework that allows for the exploration of a wider range of tasks and environments within Minecraft. BASALT competition [Shah et al., 2021] applies human evaluation to 4 open-ended tasks, and build upon that, BEDD [Milani et al., 2023] introduces a benchmark environment that systematically measures agent performance across diverse, complex scenarios
EAI Evaluation Evaluating EAI agent performance in complex environments effectively requires direct comparisons between human and AI abilities. There is a growing literature on such benchmarks, including the Animal-AI Environment framework which has been used to compare human and AI performance on cognitive tasks inspired by animal psychology, providing insights into how agents reason and generalize [Voudouris et al., 2022]. Similarly, benchmarks such as the Baby Intuitions Benchmark (BIB) assess agents’ ability to infer goals, preferences, and the actions of others, highlighting gaps in AI’s understanding of intuitive physics and social reasoning compared to humans [Gandhi et al., 2022]. These approaches underscore the challenge of developing evaluation frameworks that align with human cognitive benchmarks [Lake et al., 2016].
A.3 Societal Impacts
@@
A.4 Initial Release
Our initial release, pending the internal review process, is a 200-hour dataset. The rest of the dataset will be released over time, once they are fully vetted and participants have been given the opportunity to withdraw. To prove the existence of the full dataset, we also provide the croissant metadata for the full dataset, available to download via: https://blog.plaicraft.ai/2025/05/11/download-user-data/. Video examples from the full dataset are also given on the same blog website.
The 200-hour dataset contains the complete recordings of 3 players: Dante, Morgan and Xander. They played together as a team in the same player base throughout their time on the server, interacting constantly in most of their gameplay sessions. Thus, this subset is a perfect example of the rich social dynamics in our dataset. Dante self-identifies as a 49-year-old male player with "Regular" level experience, Morgan self-identifies as a 9-year-old male player with "Pro" level experience and Xander self-identifies as a 37-year-old female player with "Amateur" level experience.444Again, we do not guarantee the truthfulness of these information as players stay 100% anonymous during participation and information is provided completely voluntarily. Within their 200 hours of gameplay, they performed a wide range of activities, including various types of team collaborations like base construction, biomes exploration, combating each other, building iron farms, etc. Some data examples are shown in the Figure 9.
We released the croissant files for both the 200-hour dataset and the full 10,000-hour dataset, and released the raw and processed data of the 200-hour dataset at https://blog.plaicraft.ai/2025/05/11/download-user-data/.
A.5 Automatic Data Annotation
To identify subsets of highly interactive data, particularly relevant for Embodied AI, we developed a methodology leveraging audio transcribed by WhisperX [Bain et al., 2023], and the capabilities of large language models (LLMs). We first defined highly interactive data as continuous, game-related interactions between at least two players, within a specific time window. Notably, a key indicator of interactivity was the presence of rich, conversational audio. We used WhisperX-generated audio transcripts to create a dataset and identify these interactions. To preserve conversational coherence, we split the transcript into segments such that a new segment begins after a pause of at least 10 seconds between utterances. These segments represent potential areas of interactivity and Minecraft social interactions. Given the scale of the resulting dataset, manual annotation of segments was infeasible. Instead, we leveraged recent advances in large language models to perform automated annotation. A binary classification approach using a 4-shot prompt demonstrated high accuracy and efficiency on a validation subset. Based on this evaluation and overall cost considerations, we used models such as DeepSeek-V3 [DeepSeek-AI et al., 2025a] and DeepSeek-R1 [DeepSeek-AI et al., 2025b] to label the full dataset with binary annotations. Table 2 demonstrates the labeling scheme. Segments assigned specific labels were then identified as highly interactive and selected for further analysis.
@@ A.5 Automatic Data Annotation
A.6 Minecraft Configurations
To ensure stability and mod compatibility, we standardized the Minecraft version to 1.19.4 while retaining most vanilla features. Similar to MineDojo, all three worlds (the Overworld, the Nether, and the End) are accessible to players. In addition to the standard Minecraft setup, we implemented several client-side enhancements:
@@ A.6 Minecraft Configurations
We added several new worlds dedicated to minigames, isolated from the vanilla worlds. In these worlds, players can engage in a variety of minigames, such as Bedwars.
@@
We incorporated a range of third-party client-side mods and plugins to enhance performance, visuals, and playability:
Name Description Complementary Shader Pack Significantly enhances visual detail and overall graphical quality. Simple Voice Chat Proximity-based voice chat enabling real-time communication. Canary Performance optimization mod for general client-side speedups. Clumps Reduces item-entity lag by grouping dropped items into stacks. EntityCulling Skips rendering of off-screen entities to improve FPS. Fastload-Reforged Accelerates startup and world-loading times. FerriteCore Minimizes memory usage by stripping unused block states. ModernFix Addresses rendering performance issues and chunk-loading bugs. Oculus Optimizes block-update and rendering code paths for better FPS. Redirector Replaces generic implementations with faster runtime alternatives. Rubidium Speeds up core rendering operations via low-level optimizations. Saturn Bundles multiple performance tweaks for smoother gameplay. SmoothBoot Optimizes boot routines for faster game initialization. Starlight Overhauls the lighting engine for much faster chunk light updates. BetterF3 Enhanced debug overlay with customizable performance stats. Cloth-Config Provides an in-game, user-friendly configuration GUI. Entity-Model-Features Adds extra model customization hooks for entities. Entity-Texture-Features Enables per-entity texture overrides and enhancements. NotEnoughAnimation Introduces additional character animations for more realism. SkinLayer3D Renders player skin layers in true 3D for added depth. Table 3: Client-Side Mods & Plugins
@@
Various third-party server-side plugins are added:
Name Description AFK-Kick Prevents idle players by kicking after inactivity. BetterStructures Adds new, interesting auto-generated structures. Better-RTP Enhances random teleportation experience. ChatFilter Filters and blocks offensive chat content. Screaming Bedwars Implements Bedwars minigame mechanics. Chunky Preloads chunks to reduce lag spikes. CoreProtect Logs and rolls back griefing actions. DecentHolograms Displays floating holograms without dependencies. DiscordSRV Bridges Discord and Minecraft chat. EssentialsX Provides core server management commands. GravesX Spawns lootable death chests at player death. HideNametag Removes player nametags for clean visuals. InvisibleItemFrames Hides item frame visuals in-game. Multiverse-Core Manages multiple worlds on one server. PlaceholderAPI Provides a uniform placeholder system for other plugins. ProtocolLib Allows manipulation of Minecraft network packets. SuperVanish Enables admins to be invisible to players. Table 4: Server-Side Plugins
@@
A.7 Additional Participants Demographics
Our participants consist of Minecraft players across the world. To lower the overall latency players will experience, we deploy our EC2 instances in various locations across the world. Figure 11 lists out our deployment regions.
We do not hire specific participants to conduct our data collection process, as it is financially impractical, nor do we restrict anyone from participating. Instead, we focus on promoting our platform as a research project where participants can play free Minecraft with beautiful shaders on any device and make it available to everyone. We have done a series of promotions, including local poster distribution, Google Ads, YouTube and TikTok content creation. As a result, we get a constant stream of old and new participants joining our server every day. By the time of the publication, we had reached 10026 participants.

### v2

@@ 4 Evaluation
| Test Type | Spoken Instruction | Response | (ms) | (ms) | Metric |
| audio to action binding | Press A. | A pressed. | 689700 | 689850 | |
@@
| spelling ability | Please press the letter that ends the word STOP, as in S-T-O. What letter was that? | P pressed. | 115950 | 116070 | |
| object recognition | What do you have in your right hand? | Carrot. | 548097 | 548557 | |
| world-model understanding | I’m going to shoot you with an arrow, okay? And I want you to tell me how many hearts of damage are done, okay? | Two. | 510590 | 511010 | |
@@
| arithmetic | If you add up the blocks in the wool and granite pile, how many are there? | Five. | 1146213 | 1146573 | |
| spatial understanding | How many blocks away is the box? | Hundred. | 583714 | 584054 |
@@
Motivated by Animal-AI [29], which adopts testing paradigms from comparative psychology; BIB [16], which assesses machines’ abilities to reason about other agents’ intentions by observing their actions; and the taxonomy of cognitive abilities outlined in Cattell-Horn-Carroll (CHC) theory [7], we publish an evaluation suite that probes how agents perceive, understand, reason, and act within PLAICraft. The idea behind this design is intuitive: if an agent truly understands what it sees and hears, it should be able to (i) talk about it, (ii) act on it, and (iii) do so within a time‑bounded, social context.
@@
Our evaluation suite contains carefully curated prompts. Each prompt maps to one of the ten broad cognitive abilities in CHC theory [24]: fluid intelligence, quantitative knowledge, crystallized intelligence, short-term memory, visual processing, auditory processing, long-term retrieval, processing speed, decision speed, and reading/writing. Each prompt elicits a keyboard/mouse action or a concise verbal answer, enabling evaluation via exact‑match or semantic‑similarity metrics.
@@
During evaluation, the agent’s prompt consists of its entire lived life up to the current moment, including every observation and interaction so far, and is then followed by a spoken instruction that triggers the response. A correct response from the agent may be a mouse & keyboard action ("A pressed") or a brief verbal answer ("Carrot"). The test type categorizes the cognitive ability being assessed, while the metric is the measure used to judge responses’ correctness.
Table 2 selects 7 representative prompts. The prompts probe a wide range of capabilities—including visual object recognition, world-model understanding, spatial estimation, arithmetic understanding and many others. These prompts require agents to integrate visual, audio, and action in a temporally grounded context. The full list of evaluation prompts can be found in Section˜A.13.
5 Related Work
Minecraft for Machine Learning Minecraft serves as an important tool for machine learning research [20, 34] because of its open-world nature, rich dynamics, and highly customizable engine. Consequently, numerous datasets and benchmarks have been established on Minecraft [17, 21, 14, 4]. Microsoft’s MALMO platform [18] first provided a programmable interface for RL in Minecraft; building on this, MineRL [17, 25] introduced human-demonstration datasets and benchmarks for control tasks. MineDojo [14] extended this line by aggregating web-scale gameplay videos with language, while the VPT Contractor dataset [4] contributed single-player demonstrations with aligned video and keyboard/mouse actions used to train inverse-dynamics and behavioural-cloning foundation models. The BASALT competition [25] applied human evaluation to open-ended tasks, and BEDD [21] further systematized evaluation across diverse, complex Minecraft scenarios.
EAI Datasets in Other Domains Beyond Minecraft, several large-scale datasets support embodied and interactive agents in other settings. GameNGen trains on a 900M-frame corpus of VizDoom agent self-play trajectories, providing dense video–action sequences that let a diffusion model learn a real-time game engine over long horizons [28]. Mind2Web contributes crowdsourced interaction traces over 2,350 open-ended tasks on 137 real websites, using HTML/DOM state and natural-language instructions to train and evaluate generalist web agents on real UIs rather than simulators [12]. Open X-Embodiment (OXE) aggregates over 1M real-robot trajectories from 60 datasets and 22 embodiments into a unified format with egocentric RGB, proprioception, language goals, and low-level control, enabling RT-X policies that transfer across robot platforms [32]. Waymo Open provides synchronized multi-camera images, LiDAR, and HD maps for thousands of urban and suburban driving segments, and serves as a standard benchmark for perception and motion forecasting in autonomous driving [27].
We specify and compare a few of these datasets side by side with PLAICraft in Table˜3.
EAI Evaluation Evaluating EAI agent performance in complex environments effectively requires direct comparisons between human and AI abilities. There is a growing literature on such benchmarks, including the Animal-AI Environment framework which has been used to compare human and AI performance on cognitive tasks inspired by animal psychology, providing insights into how agents reason and generalize [29]. Similarly, benchmarks such as the Baby Intuitions Benchmark (BIB) assess agents’ ability to infer goals, preferences, and the actions of others, highlighting gaps in AI’s understanding of intuitive physics and social reasoning compared to humans [16]. These approaches underscore the challenge of developing evaluation frameworks that align with human cognitive benchmarks [19].
6 Discussion
| PLAICraft | MineRL [25] | MineDojo [14] | VPT Contractor [4] | Mind2Web [12] | OXE / RT-X [32] | Waymo Open [27] | ||||
| Domain | Minecraft | Minecraft | Minecraft | Minecraft | Website GUI | Robotics | Autonomous driving | |||
| Modalities | Video; hearing&speaking audio; mouse; keyboard | Video; sim state; actions | Video + text | Video; key/mouse (IDM) | HTML/DOM; instruction | RGB(+Depth); proprio; controls | RGB; LiDAR; maps/trajectories | |||
| Scale | 10,000+ h, full modalities | 500 h, full modalities | 300,000 h video; 2.2B words. | 2,000 h, full modalities | 50 h (est. from 2,350 tasks) | 2,000 h (est. from 1M+ trajectories) | 574 h | |||
| Source | 10,000+ participants | Human demonstration (# not reported) | 730K+ YouTube videos, 6K+ Wiki and 340K+ Reddit posts | 10 paid contractors from UpWork | Crowdsourced AMT demonstrations | Pooled from 60 existing lab datasets | Waymo self-driving cars in real traffic | |||
, 100 Hz; , 100 Hz; , 48 kHz | , 20 Hz; , 20 Hz | , 20 Hz; , 20 Hz | ; ; | ||||||
, 30 Hz; , 48 kHz | , 20 Hz; , | ; | , 20 Hz | ; | ; ; | , , 10 Hz; ; | |||
| Open Task | Yes | No | No | No | No | No | No | |||
| Reset | No | Episodic | Episodic | Episodic | Episodic | Episodic | Episodic | |||
| Social | Yes | No | No | No | No | No | No |
PLAICraft enables new lines of inquiry in embodied AI research. The inclusion of speech audio and the multiplayer social environments can support agents to incorporate real-time social cues, fostering collaborative play that prior single-player datasets cannot capture. The task-free, continual world and agent states also lend themselves to continual learning paradigms, where agents must develop and maintain long-horizon memory over days or weeks of game time while continually adapting to a non-stationary environment. The same long, rich trajectories, collected from the perspective of different players, can also enable systematic study of third-person learning: training agents purely from logs of other players’ behaviour rather than from their own on-policy experience. This setting, with controllable data and compute budgets, naturally supports quantitative investigation of scaling laws for embodied agents, relating training resources to control, communication, and social coordination capabilities. The vast collection of precisely aligned high-resolution videos and ambient audios—from the diverse perspectives of thousands of players in a shared world—also makes PLAICraft well-suited for world model training, 3D, and physics simulation research.
Our data collection platform is application-agnostic: because it logs raw keyboard/mouse and audio around a fixed display stream inside a controlled VM, we can easily replace Minecraft with any desktop application (e.g., web browser, productivity or creative tools) while preserving the same data schema and time-alignment guarantees. This allows us to extend beyond Minecraft agents to general computer-use agents that can operate a full computer interface (screen, audio, keyboard, mouse) at high temporal precision. The value of computer-use agents does not stop at information workflow. Remote teleoperators of warehouse and other robots are able to actuate complex robots and other cyberphysical infrastructure entirely through a standard computer interface, albeit with the added requirement of high time-precision “dexterous” keyboard and mouse actuation. Computer use agents that learn to mimic teleoperation control inputs to embodied robot agents have the potential to become de facto control policies of such robots.
Since the initial release of a subset of our dataset, it has attracted many interests from the research community. For example, Yoo et al. [33] leveraged a 50-hour subset—collected from an anonymous player (“Alex”)—to demonstrate that a video-diffusion model can be effectively trained within a continual-learning setting on a continous stream of our dataset[30], samples are shown in Figure˜9.
7 Ethical Considerations
This research ensures the safety and privacy of participants through the guidance of our university’s office of research ethics. Study procedures are reviewed routinely by the Behavioural Research Ethics Board (BREB); and have been found acceptable on ethical grounds for research involving human subjects, as well as deemed minimal risk to participants. Participants’ data is collected with their consent, or that of their guardian, and any reasonably identifiable aspects of this data are anonymized systematically.
8 Conclusion
We introduce PLAICraft, a standardized data-collection platform and large-scale multimodal dataset that provides millisecond-aligned video, speech and hearing audio, keyboard/mouse data with rich social interactions. We also introduced an evaluation suite aimed at probing the important capabilities a human-like agent should have. PLAICraft is designed to open new avenues for research in social embodied agents, continual and third-person learning, world-model, general computer-use and teleoperation-style agents.
Supplementary Material
@@ A.1 Acknowledgment
We are very grateful to Alexander Liteplo, Alex Wu, Andrew Smith, Andy Huang, Brooke Dai, Chengyuan Yao, Christopher Tardy, Daniel Crookall, David Tianyi Yin, David Yang, Edward Liang, Geo Lee, Hanson Sun, Joey Xiang, Lucas Qin, Mantaj Dhillon, Mehdi Safaee, Naveed Ghassemi, Perry Zhu, Sean Chuah, and Suzette Sun for their help in the development of the platform. We thank Hotslicer Media, Alice Xia and many others for their promotion efforts. We also appreciate the help with server moderation and community organization by David Yang, Piper, Delara, and Hiroshi (the last three are anonymous players on the server who spontaneously offered to help with moderation). In addition, we would like to send a special thanks to all the amazing players who have spent countless hours on our server, enjoying Minecraft, building a wholesome community and contributing to our research.
@@
A.2 Societal Impacts
@@
A.3 Minecraft Configurations
We retained most features of vanilla Java Minecraft. All three worlds (the Overworld, the Nether, and the End) are accessible to players. In addition to the standard Minecraft setup, we implemented several client-side enhancements:
@@ A.3 Minecraft Configurations
We built a live monitor pipeline, using WhisperX and LLM models to monitor players’ speaking and hearing audio. We then use this as metrics to determine the quality of interaction between players and reward them based on it to encourage more meaningful and high-quality interactions.
@@
We incorporated a range of third-party client-side mods and plugins to enhance performance, visuals, and playability as shown in Table˜5.
@@
Various third-party server-side plugins are added as shown in Table˜6:
| Name | Description |
| Complementary Shader Pack | Significantly enhances visual detail and overall graphical quality. |
| Simple Voice Chat | Proximity-based voice chat enabling real-time communication. |
| Canary | Performance optimization mod for general client-side speedups. |
| Clumps | Reduces item-entity lag by grouping dropped items into stacks. |
| EntityCulling | Skips rendering of off-screen entities to improve FPS. |
| Fastload-Reforged | Accelerates startup and world-loading times. |
| FerriteCore | Minimizes memory usage by stripping unused block states. |
| ModernFix | Addresses rendering performance issues and chunk-loading bugs. |
| Oculus | Optimizes block-update and rendering code paths for better FPS. |
| Redirector | Replaces generic implementations with faster runtime alternatives. |
| Rubidium | Speeds up core rendering operations via low-level optimizations. |
| Saturn | Bundles multiple performance tweaks for smoother gameplay. |
| SmoothBoot | Optimizes boot routines for faster game initialization. |
| Starlight | Overhauls the lighting engine for much faster chunk light updates. |
| BetterF3 | Enhanced debug overlay with customizable performance stats. |
| Cloth-Config | Provides an in-game, user-friendly configuration GUI. |
| Entity-Model-Features | Adds extra model customization hooks for entities. |
| Entity-Texture-Features | Enables per-entity texture overrides and enhancements. |
| NotEnoughAnimation | Introduces additional character animations for more realism. |
| SkinLayer3D | Renders player skin layers in true 3D for added depth. |
| Name | Description |
| AFK-Kick | Prevents idle players by kicking after inactivity. |
| BetterStructures | Adds new, interesting auto-generated structures. |
| Better-RTP | Enhances random teleportation experience. |
| ChatFilter | Filters and blocks offensive chat content. |
| Screaming Bedwars | Implements Bedwars minigame mechanics. |
| Chunky | Preloads chunks to reduce lag spikes. |
| CoreProtect | Logs and rolls back griefing actions. |
| DecentHolograms | Displays floating holograms without dependencies. |
| DiscordSRV | Bridges Discord and Minecraft chat. |
| EssentialsX | Provides core server management commands. |
| GravesX | Spawns lootable death chests at player death. |
| HideNametag | Removes player nametags for clean visuals. |
| InvisibleItemFrames | Hides item frame visuals in-game. |
| Multiverse-Core | Manages multiple worlds on one server. |
| PlaceholderAPI | Provides a uniform placeholder system for other plugins. |
| ProtocolLib | Allows manipulation of Minecraft network packets. |
| SuperVanish | Enables admins to be invisible to players. |
@@
A.4 Additional Participants Demographics
Our participants consist of Minecraft players across the world. To lower the overall latency players will experience, we deploy our EC2 instances in various locations across the world. Figure 11 lists out our deployment regions. Players near these areas will have a low latency when playing through our streaming. Players further away from these regions are not restricted to access our platform, but their gaming experience might suffer from lagginess due to higher latency.
We do not hire specific participants to conduct our data collection process, as it is financially impractical, nor do we restrict anyone from participating. Instead, we focus on promoting our platform as a research project where participants can play free Minecraft with beautiful shaders on any pc device and make it available to everyone. We have done a series of promotions, including local poster distribution, Google Ads, YouTube and TikTok content creation. As a result, we get a constant stream of old and new participants joining our server every day. By the time of the publication, we had reached 10026 participants.